OpenStudy (anonymous):

2 raised to the power of 5/2 - 2 raised to the power of 3/2

OpenStudy (anonymous):

When you raise a power to a power you multiply the exponents.

OpenStudy (anonymous):

well my answers are either A. 2 1/2 B. 2 C. 2 3/2 D. 2 5/3 or E. 2^2

OpenStudy (anonymous):

Well, what is (5/2-2)(3/2)

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

\[(2^{\frac{5}{2}-2})^{\frac{3}{2}}\] \[= 2^{\frac{3}{2}(\frac{5}{2}-2)}\] \[= 2^{\frac{3}{2}\frac{5}{2} - \frac{3}{2}\frac{2}{1}}\] \[= 2^{\frac{15}{4} - \frac{12}{4}}\] \[= 2^{\frac{3}{4}}\]

OpenStudy (anonymous):

Wait. I think you have a different problem.

OpenStudy (anonymous):

thats not one of the answers ,the question is 2 raised to 5 over 2 subtract 2 raised to 3 over 2?

OpenStudy (anonymous):

Yeah, you should put parens on that.. Do you mean \[2^{\frac{5}{2}} - 2^{\frac{3}{2}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Very different problem ;p

OpenStudy (anonymous):

You can factor a 1/2 from the exponents.

OpenStudy (anonymous):

Actually no you cannot.

OpenStudy (anonymous):

That's just \(\sqrt{32} - \sqrt{8}\)

OpenStudy (anonymous):

if i do 2 5/2 divided by half i get 9 and if i do 2 3/2 divided by half i get 7 so if i subtract them i obviously get 2 which is b but i dont think thats right

OpenStudy (anonymous):

\(\sqrt{32} = \sqrt{2^2*2^2*2} = 2*2\sqrt{2} = 4\sqrt{2}\)

OpenStudy (anonymous):

the answer is 2 and 3/2 but i dont no how they got that

OpenStudy (anonymous):

\(\sqrt{8} = \sqrt{2^2*2} = 2\sqrt{2}\) \[\implies \sqrt{32} - \sqrt{8} = 4\sqrt{2} - 2\sqrt{2} = 2\sqrt{2}\] So the answer should be \(2^{\frac{3}{2}}\)

OpenStudy (anonymous):

where are you getting the square roots from

OpenStudy (anonymous):

\(2^{\frac{1}{2}} = \sqrt{2}\)

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

Well think about it this way then.

OpenStudy (anonymous):

Multiplying powers of the same base you add their exponents. \(2^{5/2} \) \(= 2^{4/2} \bullet 2^{1/2}\) \(=2^2 \bullet 2^{1/2}\) \(= 4(2^{1/2})\)

OpenStudy (anonymous):

Same for \(2^{3/2} = 2(2^{1/2})\)

OpenStudy (anonymous):

So \(4(2^{1/2}) - 2(2^{1/2}) = 2(2^{1/2})\)

OpenStudy (anonymous):

\(= 2^1* 2^{1/2} = 2^{1+1/2} = 2^{3/2}\)

OpenStudy (anonymous):

so i just add 5/2 + 3/2?

OpenStudy (anonymous):

no. Watch this series on exponents.. http://www.khanacademy.org/video/exponent-properties-1?playlist=Developmental%20Math

OpenStudy (anonymous):

haha i did but i dont get subtraction... with exponents

OpenStudy (anonymous):

That's because you can't do anything special when you are adding and subtracting exponents. You can only factor out common factors. But in this case you have a common factor of 2^{1/2} that you can pull out of each term.